# Some Density Results for HyperKähler Manifolds

### v1

@@ Abstract.
Lagrangian fibrations of hyperkähler manifolds are induced by semi-ample line bundles which are isotropic with respect to the Beauville-Bogomolov-Fujiki form. For a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ of positive dimension, we prove that the set of points in $S$, for which there is an isotropic class in the Picard lattice of the corresponding hyperkähler manifold represented as a fiber over that point, is analytically dense in $S$. We also prove the expected openness and density of the locus of polarised hyperkähler manifolds that admit a nef algebraic isotropic line bundle.
@@ Introduction
Since any hyperkähler manifold $X$ comes equipped with a holomorphic symplectic form $\sigma$, Lagrangian subvarieties of $X$, i.e., subvarieties $Z\subset X$ of dimension $n=\frac{1}{2}\dim X$ such that $\sigma|_{Z^{\rm reg}}=0$, are interesting to study. It is known [Mat99] that any morphism $f\colon X\to B$ where $0<\dim B<2n$, is in fact a Lagrangian fibration, meaning the general fiber of $f$ is a Lagrangian subvariety. The general fibers are in fact abelian varieties, as proved by Voisin (see [Cam06, Prop. 2.1]). Moreover, whenever $B$ is smooth, it is in fact isomorphic to $\mathbb{P}^{n}$ [Hwa08]. Hence, such a Lagrangian fibration is induced by a semi-ample line bundle $\mathcal{L}$ on $X$ with $c_{1}(\mathcal{L})^{n+1}=0$.
The second cohomology $H^{2}(X,{\mathbb{Z}})$ comes equipped with a unique non-degenerate quadratic form $q_{X}$, the Beaville–Bogomolov–Fujiki form, which controls a great deal of geometric information about the hyperkähler manifold $X$. It has signature $(3,b_{2}-3),$ and there exists a non-zero constant $c_{X}\in\mathbb{Q}$ satisfying $\alpha^{2n}=c_{X}\cdot q_{X}(\alpha)^{n}$. In particular, admitting a semi-ample line bundle $\mathcal{L}$ with $c_{1}(\mathcal{L})^{n+1}=0$ is equivalent to $q_{X}(\mathcal{L})=0.$ Therefore, the first step in studying Lagrangian fibrations is the study of isotropic classes.
@@
Let ${\mathcal{X}}\rightarrow S$ be a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ with $\dim S>0.$ Then the set
$\{t\in S\mid\exists\alpha\in H^{1,1}({\mathcal{X}}_{t},\mathbb{Z}),q_{X}(\alpha)=0\}$
is dense in $S$ in the analytic topology.
The case when $S=\operatorname{Def}(X)$, the whole deformation space of $X$, is already known. For hyperkähler manifolds, there is a global Torelli theorem which allows one to prove such density statements directly on the period domain (i.e., the type IV symmetric domain associated to $(H^{2}(X,\mathbb{Z}),q_{X})$; see §1 for details). On the period domain, these types of density results follow from ergodic properties [Huy16, Prop. 7.1.3]. However, the question of density in an arbitrary family demands different methods. Our argument is Hodge theoretic and hence applies more generally to any non-isotrivial family of K3-type Hodge structures (see Theorem 1.2 for a precise statement). We follow [Voi07, §5.3.4].
One should compare Theorem 1.2 to the classical density result for Noether–Lefschetz loci, which, in the set-up of the theorem, states that the locus of hyperkähler manifolds where the Picard rank jumps is analytically dense in the base. The result is attributed to Oguiso [Ogu03, Theorem 1.1] and the key ingredient (also a key ingredient in our proof) is the criterion for the density of the Hodge classes, due to Green [Voi07, Proposition 5.20].
In the pursuit of Lagrangian fibrations, algebraic isotropic111isotropic = square zero classes as above are in a way insufficient; a hyperkähler manifold $X$ may have such an algebraic isotropic class $\alpha$, while none of the elements in its orbit under birational self-maps of $X$ are nef (see [Sac23, Thm. 4.1] for an example). Nefness is a necessary condition for such an $\alpha$ to induce a Lagrangian fibration. The hyperkähler SYZ conjecture postulates that it is in fact sufficient. The conjecture was first formulated by Strominger, Yau and Zaslow in [SYZ96] in the context of mirror symmetry of Calabi–Yau pairs. It has been reformulated for hyperkähler manifolds by Huybrechts, Hasset–Tschinkel, Tyurin, Bogomolov, and Sawon.
@@
    A hyperkähler manifold $X$ admits a rational Lagrangian fibration (i.e., it is birational to a hyperkähler manifold which admits a Lagrangian fibration) if and only if it admits an isotropic non-trivial line bundle $L$.
Note that, according to [KL22, Lemma 2.18], for a fixed hyperkähler manifold $X$, then (1) is satisfied for all hyperkählers deformation equivalent to $X$ if and only if (2) is satisfied for all hyperkählers deformation to $X$. Therefore, one may focus on isotropic nef classes. To this end, Matsushita and Kamenova–Verbitsky studied the locus $\operatorname{Def}(X,L)_{\operatorname{reg}}$ of hyperkähler manifolds which admit a regular Lagrangian fibration in the deformation space $\operatorname{Def}(X,L),$ where $L$ is an algebraic isotropic line bundle. They showed:
@@ 1. Density of isotropic classes
Throughout, we denote by $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ a polarised variation of Hodge structures (PVHS) over a connected complex manifold $S$. Here ${\mathbb{V}}$ is a local system of finite dimensional ${\mathbb{Q}}$-vector spaces with a Hodge filtration ${\mathcal{F}}^{\bullet}$ of the holomorphic bundle ${\mathcal{V}}:={\mathbb{V}}\otimes{\mathcal{O}}_{S}$, and $\nabla\colon{\mathcal{V}}\rightarrow{\mathcal{V}}\otimes\Omega_{S}^{1}$ is the Gauss-Manin connection. The polarisation $q$ is a flat bilinear form ${\mathbb{V}}\otimes{\mathbb{V}}\rightarrow{\mathbb{V}}$ whose restriction to each fiber $V_{s}$ of ${\mathbb{V}}$ at a point $s\in S$ is a polarisation of the Hodge structure on $V_{s}$.
@@
A finite dimensional vector space $V_{\mathbb{Q}}$ with a quadratic form $q$ is of K3-type if the signature of $q$ is $(3,\dim V_{\mathbb{R}}-3)$ with $\dim V_{\mathbb{Q}}\geq 4$. We say that a PVHS $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ is of $K3$-type if it is a variation of polarised weight 2 Hodge structures with $\operatorname{rank}\mathcal{F}^{2}=1$ and each fiber is a K3-type vector space.
@@
Let $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ be a PVHS of K3-type with a trivialisation $\varphi\colon\mathbb{V}\overset{\sim}{\to}V_{\mathbb{Q}}\otimes\mathbb{Q}_{S}$. We have the holomorphic period map:
@@
defined by $t\mapsto\varphi_{t}(\mathcal{F}^{2}_{t})$. We say that a PVHS of K3-type is non-isotrivial if the holomorphic period map $\mathcal{P}$ is non-constant.
@@
We let $Q:=\{x\mid q(x)=0\}\subset V_{\mathbb{C}}$ be the quadric cone defined by the polarisation; we will assume that $Q$ has a nonzero rational point. Note that, if $\operatorname{rank}({\mathbb{V}})\geq 5$, this is automatically satisfied by Meyer’s theorem.
@@
Let $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ be a non-isotrivial PVHS of $K3$ type with trivialisation $\varphi\colon{\mathbb{V}}\overset{\sim}{\to}V_{\mathbb{Q}}\otimes{\mathbb{Q}}_{S}$ as above. Assume that the quadric cone $Q$ has a nonzero rational point. Then the union
$S_{\rm isot}:=\bigcup_{\begin{subarray}{c}\alpha\in V_{\mathbb{Q}}\\ q(\alpha)=0\end{subarray}}\alpha^{\perp}(S)$
@@
and denote $p\colon Q^{1}\rightarrow Q$ the restriction of the projection to the first factor. By Lemma 1.3 below, the map $p$ is open around some point $(\lambda,s)\in Q^{1}$. Let $\pi\colon Q^{1}\to S$ be the restriction of the second projection. For any $s\in S$, there is an open neighbourhood $U\ni s$ such that $\pi^{-1}(U)\subset Q^{1}$ contains an open neighbourhood $V$ of $(\lambda,s)$ surjecting onto $U$ where $p|_{V}\colon V\to Q$ is open.
Note that the induced map on the real points $p_{\mathbb{R}}\colon V(\mathbb{R})\to Q(\mathbb{R})$ is also open. The sets $Q({\mathbb{Q}})$ and $Q({\mathbb{R}})$ are non-empty by assumption. Therefore, since $Q({\mathbb{Q}})\subset Q({\mathbb{R}})$ is dense, it follows that the subset $\{(\alpha,s)\mid\alpha\in\mathcal{F}^{1}_{s},q(\alpha)=0\}\cap V\subset V$ is dense. Therefore its image $\bigcup_{q(\alpha)=0}\alpha^{\perp}(U)$ in $U$ is also dense. ∎
@@
Denote by $\widetilde{p}\colon\mathcal{F}^{1}\longrightarrow V_{\mathbb{C}}$ the projection to the first factor. By [Voi07, Lemma 5.21],
the map $\widetilde{p}$ is a submersion at $(\lambda,s)\in\mathcal{F}^{1}$ if the map
@@
Since the PVHS $(S,{\mathbb{V}},\mathcal{F}^{\bullet},\nabla,q)$ is not isotrivial, this differential is not zero and hence there exists $v\in T_{S,s}$ such that $d\mathcal{P}(v)\colon V_{s}^{1,1}\to V^{0,2}_{s}$ is non-trivial, hence surjective. Therefore there exists $\lambda\in V_{s}^{1,1}$ such that ¯$\overline{\nabla}(\lambda)$ is surjective.
For a fixed $s$, the quadric $Q^{1}_{s}\cap V^{1,1}_{s}\subset V^{1,1}_{s}$ is not contained in a hyperplane by the definition of a polarisation. Since the dimension of $V_{s}^{1,1}$ is at least 2, $Q^{1}_{s}\cap V^{1,1}_{s}$ cannot be contained in the linear subspace $\operatorname{Ker}(d\mathcal{P}(v))\subset V_{s}^{1,1}$. Therefore, the map $\widetilde{p}$ is a submersion at $\lambda\in Q^{1}_{s}$.
@@
[Theorem 1] Let ${\mathcal{X}}\rightarrow S$ be a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ with $\dim S>0.$ Then the union
$S_{\rm isot}:=\bigcup_{q(\alpha)=0}\alpha^{\perp}(S)$
is dense in $S$ in the analytic topology.
@@
One could also apply Theorem 1.2 to the Fano variety $F(X)$ of lines associated to a cubic 4-fold $X$. The Fano variety $F(X)$ is a projective hyperkähler manifold of K3${}^{[2]}$-type. By [Huy23, Corollary 5.24], if $F(X)$ admits a rational Lagrangian fibration, then the cubic fourfold $X$ lies on a special divisor $\mathscr{C}_{d}$ of the moduli of cubic fourfolds, indexed by an integer $d$ where $d/2$ is a perfect square. Said differently, Theorem 1.2 implies that any non-isotrivial family of cubic fourfolds contains a dense subset of points from $\mathscr{C}_{d}$ where $d/2$ is a perfect square.

### v3

@@ Abstract.
Lagrangian fibrations of hyperkähler manifolds are induced by semi-ample line bundles which are isotropic with respect to the Beauville–Bogomolov–Fujiki form. For a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ of positive dimension, we prove that the set of points in $S$, for which there is an isotropic class in the Picard lattice of the corresponding hyperkähler manifold represented as a fiber over that point, is analytically dense in $S$. We also prove the expected openness and density of the locus of polarised hyperkähler manifolds that admit a nef algebraic isotropic line bundle.
@@ Introduction
Since any hyperkähler manifold $X$ comes equipped with a holomorphic symplectic form $\sigma$, Lagrangian subvarieties of $X$, i.e., subvarieties $Z\subset X$ of dimension $n=\frac{1}{2}\dim X$ such that $\sigma|_{Z^{\rm reg}}=0$, are interesting to study. It is known [Mat99] that any morphism $f\colon X\to B$ where $B$ is a normal complex variety with $0<\dim B<2n$, is in fact a Lagrangian fibration, meaning a surjective morphism with connected fibers where the general fiber is a Lagrangian subvariety. The general fibers are in fact abelian varieties, as proved by Voisin (see [Cam06, Prop. 2.1]). If follows from [Mat16, Remark 1.4] that the base $B$ is projective. Moreover, whenever $B$ is smooth, it is in fact isomorphic to $\mathbb{P}^{n}$ [Hwa08, GL14]. A Lagrangian fibration with base $B\cong\mathbb{P}^{n}$ is induced by a line bundle $\mathcal{L}$ on $X$ with $c_{1}(\mathcal{L})^{n+1}=0$.
The second cohomology $H^{2}(X,{\mathbb{Z}})$ comes equipped with a unique non-degenerate quadratic form $q_{X}$, the Beaville–Bogomolov–Fujiki form, which controls a great deal of geometric information about the hyperkähler manifold $X$. It has signature $(3,b_{2}-3),$ and there exists a non-zero constant $c_{X}\in\mathbb{Q}$ satisfying $\alpha^{2n}=c_{X}\cdot q_{X}(\alpha)^{n}$. In particular, admitting a line bundle $\mathcal{L}$ with $c_{1}(\mathcal{L})^{n+1}=0$ is equivalent to $q_{X}(\mathcal{L})=0.$ Therefore, the first step in studying Lagrangian fibrations is the study of isotropic classes.
@@
Let ${\mathcal{X}}\rightarrow S$ be a non-isotrivial family of hyperkähler manifolds with $b_{2}(X_{t})\geq 5$ over a complex manifold $S$ with $\dim S>0.$ Then the set
$\{t\in S\mid\exists\alpha\in H^{1,1}({\mathcal{X}}_{t},\mathbb{Z}),\alpha\neq 0,q_{X}(\alpha)=0\}$
is either empty or dense in $S$ in the analytic topology.
The case when $S=\operatorname{Def}(X)$, the whole deformation space of $X$, is already known. For hyperkähler manifolds, there is a global Torelli theorem which allows one to prove such density statements directly on the period domain (i.e., the type IV symmetric domain associated to $(H^{2}(X,\mathbb{Z}),q_{X})$; see §1 for details). On the period domain, these types of density results follow from ergodic properties [Huy16, Prop. 7.1.3]. For certain moduli spaces of lattice polarised hyperkähler manifolds, Mongardi and Pacienza establish similar density results in [MP23, Theorem 3.15]. However, the question of density in an arbitrary family demands different methods. Our argument is Hodge theoretic and hence applies more generally to any non-isotrivial family of K3-type Hodge structures (see Theorem 1.2 for a precise statement).
One should compare Theorem 1.2 to the classical density result for Noether–Lefschetz loci, which, in the set-up of the theorem, states that the locus of hyperkähler manifolds where the Picard rank jumps is analytically dense in the base. This result is attributed to Oguiso [Ogu03, Theorem 1.1] and the key ingredient (also a key ingredient in our proof) is the criterion for the density of the Hodge classes, due to Green [Voi07, Proposition 5.20]. For certain families of polarized hyperkähler manifolds, the density of the loci of hyperkählers with algebraic isotropic classes follows from equidistribution results of Tayou [Tay20], and Tayou and Tholozan [TT23].
Nefness is a necessary condition for an algebraic isotropic class $\alpha$ to induce a Lagrangian fibration. The hyperkähler SYZ conjecture postulates that it is in fact sufficient. The conjecture was first formulated by Strominger, Yau and Zaslow in [SYZ96] in the context of mirror symmetry of Calabi–Yau pairs. It has been reformulated for hyperkähler manifolds by Huybrechts, Hasset–Tschinkel, Tyurin, Bogomolov, and Sawon.
@@
    A hyperkähler manifold $X$ admits a rational Lagrangian fibration (i.e., it is birational to a hyperkähler manifold which admits a Lagrangian fibration) if and only if it admits a non-trivial isotropic line bundle $L$.
Note that in (2) the isotropic line bundle $L$ is not required to be nef. An explicit example of a hyperkähler manifold with a strictly rational Lagrangian fibration satisfying (2) was found in [Sac23, Thm. 4.1], by performing a Mukai flop along the image of a section of a regular Lagrangian fibration. Further, according to [KL22, Lemma 2.18], for a fixed hyperkähler manifold $X$, then (1) is satisfied for all hyperkählers deformation equivalent to $X$ if and only if (2) is satisfied for all hyperkählers deformation to $X$. Therefore, one may focus on isotropic nef classes. To this end, Matsushita and Kamenova–Verbitsky studied the locus $\operatorname{Def}(X,L)_{\operatorname{reg}}$ of hyperkähler manifolds which admit a regular Lagrangian fibration in the deformation space $\operatorname{Def}(X,L),$ where $L$ is an algebraic isotropic line bundle. They showed:
@@ 1. Density of isotropic classes
Throughout, we denote by $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ a variation of Hodge structures (VHS), with a bilinear form $q$, over a connected complex manifold $S$. Here ${\mathbb{V}}$ is a local system of finite dimensional ${\mathbb{Q}}$-vector spaces with a Hodge filtration ${\mathcal{F}}^{\bullet}$ of the holomorphic bundle ${\mathcal{V}}:={\mathbb{V}}\otimes{\mathcal{O}}_{S}$, and $\nabla\colon{\mathcal{V}}\rightarrow{\mathcal{V}}\otimes\Omega_{S}^{1}$ is the Gauss–Manin connection. The form $q$ is a flat bilinear form ${\mathbb{V}}\otimes{\mathbb{V}}\rightarrow{\mathbb{Q}}_{S}$ whose restriction to each fiber $V_{s}$ of ${\mathbb{V}}$ at a point $s\in S$ is non-degenerate on $V_{s}$.
@@
A finite dimensional vector space $V_{\mathbb{Q}}$ with a quadratic form $q$ is of K3-type if the signature of $q$ is $(3,\dim V_{\mathbb{R}}-3)$ with $\dim V_{\mathbb{Q}}\geq 4$. We say that a VHS $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ is of $K3$-type if it is a variation of weight 2 Hodge structures with $\operatorname{rank}\mathcal{F}^{2}=1$ and each fiber is a K3-type vector space.
@@
Let $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ be a VHS of K3-type with a trivialisation $\varphi\colon\mathbb{V}\overset{\sim}{\to}V_{\mathbb{Q}}\otimes\mathbb{Q}_{S}$. We have the holomorphic period map:
@@
defined by $t\mapsto\varphi_{t}(\mathcal{F}^{2}_{t})$. We say that a VHS of K3-type is non-isotrivial if the holomorphic period map $\mathcal{P}$ is non-constant.
@@
We let $Q:=\{x\mid q(x)=0\}\subset V_{\mathbb{C}}$ be the quadric cone defined by the form $q$; we will assume that $Q$ has a nonzero rational point. Note that, if $\operatorname{rank}({\mathbb{V}})\geq 5$, this is automatically satisfied by Meyer’s theorem.
@@
Let $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ be a non-isotrivial VHS of $K3$ type with trivialisation $\varphi\colon{\mathbb{V}}\overset{\sim}{\to}V_{\mathbb{Q}}\otimes{\mathbb{Q}}_{S}$ as above. Assume that the quadric cone $Q$ has a nonzero rational point. Then the union
$S_{\rm isot}:=\bigcup_{\begin{subarray}{c}\alpha\in V_{\mathbb{Q}}\\ \alpha\neq 0\\ q(\alpha)=0\end{subarray}}\alpha^{\perp}(S)$
@@
and denote $p\colon Q^{1}\rightarrow Q$ the restriction of the projection to the first factor. By Lemma 1.3 below, for any $s\in S$, the map $p$ is open around some point $(\lambda,s)\in Q^{1}$. Let $\pi\colon Q^{1}\to S$ be the restriction of the second projection. For any $s\in S$, there is an open neighbourhood $U\ni s$ such that $\pi^{-1}(U)\subset Q^{1}$ contains an open neighbourhood $W$ of $(\lambda,s)$ surjecting onto $U$ where $p|_{W}\colon W\to Q$ is open.
Note that the induced map on the real points $p_{\mathbb{R}}\colon W(\mathbb{R})\to Q(\mathbb{R})$ is also open. The sets $Q({\mathbb{Q}})$ and $Q({\mathbb{R}})$ are non-empty by assumption. Therefore, since $Q({\mathbb{Q}})\subset Q({\mathbb{R}})$ is dense, it follows that the subset $\{(\alpha,s)\mid\alpha\in\mathcal{F}^{1}_{s},\alpha\neq 0,q(\alpha)=0\}\cap W\subset W$ is dense. Therefore its image $\bigcup_{\begin{subarray}{c}\alpha\neq 0,\\ q(\alpha)=0\end{subarray}}\alpha^{\perp}(U)$ in $U$ is also dense. ∎
@@
Denote by $\widetilde{p}\colon\mathcal{F}^{1}\longrightarrow V_{\mathbb{C}}$ the projection to the first factor. By [Voi07, Lemma 5.21], the map $\widetilde{p}$ is a submersion at $(\lambda,s)\in\mathcal{F}^{1}$ if the map
@@
Since the VHS $(S,{\mathbb{V}},\mathcal{F}^{\bullet},\nabla,q)$ is not isotrivial, this differential is not zero and hence there exists $v\in T_{S,s}$ such that $d\mathcal{P}(v)\colon V_{s}^{1,1}\to V^{0,2}_{s}$ is non-trivial, hence surjective. Therefore there exists $\lambda\in V_{s}^{1,1}$ such that ¯$\overline{\nabla}(\lambda)$ is surjective.
For a fixed $s$, the quadric $Q^{1}_{s}\cap V^{1,1}_{s}\subset V^{1,1}_{s}$ is not contained in a hyperplane by the non-degeneracy of the quadratic form. Since the dimension of $V_{s}^{1,1}$ is at least 2, $Q^{1}_{s}\cap V^{1,1}_{s}$ cannot be contained in the linear subspace $\operatorname{Ker}(d\mathcal{P}(v))\subset V_{s}^{1,1}$. Therefore, the map $\widetilde{p}$ is a submersion at $\lambda\in Q^{1}_{s}$.
@@
(= Theorem 1) Let ${\mathcal{X}}\rightarrow S$ be a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ with $\dim S>0$ and $b_{2}(X_{t})\geq 5$. Then the union
$S_{\rm isot}:=\bigcup_{\begin{subarray}{c}\alpha\neq 0,\\ q(\alpha)=0\end{subarray}}\alpha^{\perp}(S)$
is either empty or dense in $S$ in the analytic topology.
@@
One could also apply Theorem 1.2 to the Fano variety $F(X)$ of lines associated to a cubic 4-fold $X$. The Fano variety $F(X)$ is a projective hyperkähler manifold of K3[2]-type. By [Huy23, Corollary 5.24], if $F(X)$ admits a rational Lagrangian fibration, then the cubic fourfold $X$ lies on a special divisor $\mathscr{C}_{d}$ of the moduli of cubic fourfolds, indexed by an integer $d$ where $d/2$ is a perfect square. Said differently, Theorem 1.2 implies that any non-isotrivial family of cubic fourfolds contains a dense subset of points from $\mathscr{C}_{d}$ where $d/2$ is a perfect square.
@@
Remark 1.6.
Notice that the period domain $\Omega$ contains smooth compact Riemann surfaces of any genus, however it does not contain compact complex subspaces of dimension greater than one (see the discussion in [GS23, Sections 5.6 and 5.7]). Therefore, the image $\mathcal{P}(S)$ of the complex manifold $S$ inside the period domain $\Omega$ is either non-compact of any dimension, or compact of dimension one.